At an amusement park, the cost of 3 adult tickets and 4 children's tickets is $126.50. Another customer paid $120.50 for 2 adult tickets and 5 children's tickets. What is the price of each kind of ticket?

Answer by checkley79(3341)

3A+4C=126.50 MULTIPLY BY -2.
2A+5C=120.50 MULTIPLY BY 3 & ADD.
-6A-8C=-253.00
6A+15C=361,50
--------------------
7C=108.50
C=108.50/7
C=$15.50 PRICE OF A CHILDREN'S TICKET.
3A+4*15.50=126.50
3A+62=126,50
3A=126.50-62
3A=64.50
A=64.50/3
A=$21.50 PRICE OF AN ADULT TICKET.
PROOF:
2*21.50+5*15.50=120.50
43+77.50=120.50
120.50=120.50